FEL Upgrade Project Weekly Brief –June 10-14, 2002




We warmed up the injector cryounit this week so that we could replace the rf windows for the higher powers required for the 10 kW upgrade, replace a leaky valve, and check the alignment prior to installing new gun and injector hardware.


The draft report from the May 1-3 semiannual review of the FEL Upgrade project was reviewed by the project team and sent back to the ONR contract monitor for distribution. We thank the panel for their constructive and fair review of our progress.

The project financial reports for the month of May were received and we began the summary process so that the May monthly report could be completed and distributed next week.

Next week two separate meetings of JTO Technical Working Groups are scheduled which will require JLab participation: the lethality working group will meet on the 18th at the Dahlgren NSWC and the FEL working group will meet on the 21st at the University of Maryland.

On Monday, June 10, F.Dylla met with the management of the Wellman Laboratory of Photomedicine at Massachusetts General Hospital in Boston to continue our discussions for collabortive efforts on FEL applications to photobiology and photomedicine.

WBS 4 (Injector):

The shield door was implanted successfully. A HEPA filter unit and new curtains have been installed around the W&M implant system creating a much cleaner area. The cables on the system were replaced with shielded cables to reduce noise. The ball cathode requires additional polishing and is being worked. The SF6 tank modification drawings are in review.

Gun HVPS - The SF6 tank is due next week. The Conditioning Resistor is having problems due to the firing temperature profile. If the resistive material remains above 650°C for too long, glass in the resistive compound encapsulates the conductive particles and the resistor becomes an open. We are working with the heat treatment vendor to get the heat-up and cool-down times to be a few minutes between 650°C and 850°C. A lower resistivity material will be tried next by screening it on to the ceramic cylinder.

WBS 6 (RF):

Quarter Klystrons - The 2nd and 3rd klystron carts were modified by the Machine Shop this week and returned. The solenoids and klystrons will be mounted next week.

Quarter HVPS - The wiring for the new PLC was completed this week for unit 3. Armor flex is being ordered to contain the HV wiring to the klystron carts. Testing of the PLC's is expected to start next week. The system drawings are being kept up to date with all changes and modifications. A precision rectifier circuit board was designed, built, calibrated, and installed. A printed circuit board is in Drafting and is being laid out for the production units.

WBS 8 (Instrumentation):

The last magnet trim rack feeds have been cabled into the junction box downstairs in the vault. A 3-phase transformer and breaker panel for the magnet power supplies has been installed in the west end of the gallery. This will simplify the AC power connections to the new magnet supplies and enhance safety by providing a single point to lock and tag the new power supplies.

The final cabling is being done for the girders installed in the tunnel. All the light rings for the

beam viewers have been stuffed and tested. Camera assemblies for MQX3F04, MQX3F05, MQX3F06 have been installed.

Work continued on the fabrication for the Beam Loss Monitors. The parts for the Photo Multiplier Tube (PMT) are all in and are being assembled by one of our students. The new Canbus Interface board for the beam viewer control chassis upgrade has been completed and is at the board house with a delivery date of 10 days.

Six of the BLM rear board prototypes (F0150) have been sent out for fabrication to be received on a 5-day turn. Front panels for the boards have been ordered as well and all the required parts are on order. Final adjustments to the BLM front card (F0151) schematic have been made and the package is now back with EECAD. Prototypes are planed to go out for fabrication in early July. Improvements continue on the inventory control system. An update to the posted site is likely to occur early next week.

Additional improvements have been made to the ion implantation system. A remote control box is being discussed such that the metal rack door could be kept closed during processing to help shield electrical noise from the Pulse Forming Network (PFN).

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland, CA, we have a nagging cosmetic problem with the surface of the coils after potting. After removal from the potting mold, within an hour, welt-like ridges start to form. They are .005 high, with white, flake-like spots generate out of the originally flat, clear surfaces. The ridges and spots nominally form above the space between conductors, mostly in the flats where the bundle is squeezed to make the coil fit in the coil gap.

Our epoxy consultant, Jeffrey Hubrig of Innovation Services tracked down a case where similar welts in the epoxy outer surface along the zone between conductors was solved. The manufacturer out gassed the material at higher temperature for longer times.

We are theorizing that the Mylar tape requires out gassing at a higher temperature than in the past because of possible reformulation of its adhesive. On the next coil (#12) yesterday, MEI bake out at 270°F(rather than 130) for 16 hours (rather than 4) and degassed for 4 to 6 hours (rather than 1.5) Hopefully, this will fix the problem as in the 1984 case. Are coils degraded with this surface imperfection? Well, they are not pristine looking coils. But the primary insulation, the mylar, is undisturbed. The glass tape and epoxy are to hold the coil together and provide secondary insulation. This material is slightly disturbed by rework. MEI is authorized to and will rework the surfaces with sanding within the gap zone to eliminate the welts. They will then coat the surface with a Miller and Stevenson clear conformal spray. They follow this with a local hi-pot test using a brillo pad. All reworked coils have passed hipot so far.

• All four core plate sets for the IR machine passed gap inspection, having uniformity variation of no greater than .0005 inch (1/4 human hair) The cores are now at MEI going through inspection. The first unit will go through a pre assembly of al its parts to verify bolt and part fit-up.

• MEI also performed a successful prototyping of their Purcell Gap Shim glue down procedure within a freshly constructed clean room. The shims were uniformly glued to the base plate within .0005. They will glue the first shims on Monday to a real core.

• Cores for the Second batch of 6 UV Magnets remain in final machining.

Injector Dipoles (DU/DV)

• At Master Machine, GV cores are complete except for painting. GU cores are now in final machining.

• At Wang NMR, GU/GV coil potting started this week.

Arc 180 Degree Dipoles (GY)

• The assembled 2nd Core passed its gap check at Bosma. I quote from my acceptance memo: "If I look at it the way the wide beam of electrons sees this gap and its field, that is at any one radial set of measurements, the uniformity among measurements is within .0002 inches.

• At Wang NMR, the second GY Coil (of four) was completed.

• GY Coil Potting fixture is in fabrication.

• GG Coils are being potted.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) continued manufacturing the GQ-GX Cores.

• GQ upper, lower cores remain in final machining.

• GX cores are through heat treat.

• Their plan for first article GQ is now July 9.

• We funded their price change to add shims to the magnet pole faces for maximum


3 inch quad (QX)

• We completed measuring all QX quads for both the IR and the UV machine.

Trim Quad (QT)

• Milhous Control of Virginia plans to ship three magnets today.

• At last a potential solution to the four required vertical correctors in the arcs. (GCs) Tom Hiatt's analysis shows that the Trim Quad (QT) is a perfect base for an accurate vertical dipole corrector if the main top and bottom coils are run at 1.6 A and we fill in the vacant corners of the Panofsky coils of the quad with separate coils run at about the same current density and in the proper vertical corrector dipole sense.

• The only downside is that to power the upper/lower coils in this configuration, where one has subtracted current (from the quadrupole current) and the other has current addition, we need two power supplies per corrector magnet that have to be ganged to run in parallel at the same current,

Sextupole (SF)

• DULY Research supplied us with check prints of the magnet. We also made a last minute switch in the pinning/bolting configuration that went to single 5/8 bolts and 3/8 pins that have larger distances between so that the structure is not compromised.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The procurement documents for the ARC Chambers were completed and the drawings put out for electronic distribution via our web procurement system.

• The X Chamber remains in fabrication in the Shop.

• The drawings for the recirculation region in front of the Cryomodule will be signed off on Monday.

• Design of three of the four chambers for the optical chicane region is now complete.

• Design of the regions around the wiggler continues.



WBS 11 (Optics):